# Belinda Z. Li    Zifan Carl Guo    Jacob Andreas

### v2

@@ Appendix C Full Probing Results
The mean probability the probe put on the correct answer is plotted in Figure 9. We find that generally speaking, both AA and PAA linearly encode states of exponentially longer sequences as they go down the layers. We find evidence that the PAA models use their intermediate layers to compute parity in parallel: at around the second residual layer, PAA models place $\frac{1}{3}$ probability on the correct answer (there are three actions of each parity in $S_{3}$).
C.1 Examples of Associative Algorithm Representations that Do or Do Not Linearly Encode Parity
@@
Figure 10: Example activations from an AA model that does not linearly encodes parity at the final layer, projected on three principal with a total explained variance of 41.8%. Blue points have even parity, while orange points have odd parity.
Figure 11: Example activations from an AA model that does linearly encodes parity at the final layer. Blue points have even parity, while orange points have odd parity. (Left) State and state parity probe signatures of this model. (Right) projection of hidden representations onto three components with a total explained variance of 49.9%.
@@ D.1 $S_{3}$
We visualize the linear decomposition of the last-layer or penultimate-layer representations across various different PAA models. We find the triangular prism shape similar to Figure 4 in all of them, but there was no consistency in which states were paired to form the clusters.
One interpretation is that PAA models may be learning various presentations of $S_{3}$, with each clustering configuration corresponding to a different presentation. Generally speaking, $S_{n}$ can be generated by a 2-cycle and a $n$-cycle: any permutation of $S_{n}$ can be created by composing these two permutations. For example, $S_{3}$ can be generated by the 2-cycle $1\leftrightarrow 2$ and 3-cycle $1\to 2\to 3\to 1$, which corresponds to the clustering $\{(123,213),(312,132),(231,321)\}$: the states within the cluster can be transformed into each other by applying $1\leftrightarrow 2$, while states between clusters are related to each other by $1\to 2\to 3\to 1$. PAA models that cluster according to this pattern may have learned these generators.
@@ D.2 $S_{5}$
What happens in models that learn PAA in $S_{5}$? We visualize the penultimate-layer representation of a Pythia-160M model that learned PAA on $S_{5}$ in 3D space, with parity along one axis and two orthogonal directions along the other two. We find 4 clearly separated clusters, corresponding to the position of 1 in the state (states having 1 in position 4 and 1 in position 5 are clustered together).
@@
Appendix E Full Attention Heads Analysis
E.1 Formalizing Parity Heads
@@
Definition E.1.
Let $\alpha_{i,\ell}^{(H)}(x)$ be the attention weight of the $H$th attention head in layer $\ell$ at position $i$ for input $x$ where $x$ is the list of actions $[a_{1}\ldots a_{t}]$.
@@
Here, $[a_{1}\ldots a_{i}]$ denotes the first $i$ elements of the sequence $x$, and $\text{CI}_{\ell,H,\text{odd}}$ refers to the confidence interval around the average attention weights on odd-parity tokens. We use sequence lengths of up to $L=80$ for $S_{3}$ and $L=50$ for $S_{5}$.
@@
E.2 AA Attention Patterns
Figure 14: Attention patterns in AA models form a tree-like pattern, with tokens in successive layers attending to larger windows of downstream tokens. This is in line with how we expect the associative algorithm to function. We plot only the most salient attention weights by pruning edges for which attention weight is $<0.95$, the target token is not in the top-3 attended-to tokens from the source token, or the source token is not in the top-10 attended-from tokens for the target token. We expect that the attention patterns visualized here do not form a single clean tree, but the superimposition of multiple trees.
@@ E.2 AA Attention Patterns
    Edges between two nodes $(t_{1},l-1)$ and $(t_{2},l)$ if position $t_{2}$ at layer $l$ attends to position $t$ at layer $l-1$. We define “attends to” as follows:
    Let $\alpha_{t_{1}\rightarrow t_{2}}^{(l)}$ denote the maximum attention weight (across all attention heads at layer $l$) from position $t_{1}$ at layer $l-1$ to position $t_{2}$ at layer $l$.
@@
      $\alpha_{t_{1}\rightarrow t_{2}}^{(l)}\in\text{top-10}(\{\alpha_{t_{1}\rightarrow j}^{(l)}:k\in[1,n]\})$: $t_{2}$ is among the top-10 attended-from tokens for $t_{1}$ at layer $l-1$
@@
Note that the tree is not entirely clean: there are redundant edges and edges that cross each other: we suspect that the subgraph we’ve picked up on is not a single tree, but rather the superimposition of multiple trees, each potentially contributing to not just the prediction for the final token, but also the predictions of the previous tokens.
Appendix F How are these algorithms learned over the course of training?
@@
Appendix G Additional Factors Influencing Learned Algorithm
G.1 Model Size
We have investigated whether model size influences which algorithm the models learn to implement. As shown in Figure 15, model size empirically does not seem to have much effect on the choice of learned algorithm, with model architecture and initialization playing a much bigger effect. Some notable differences between the GPT-2 and Pythia architecture are the use of rotary embeddings, parallelized attention and feedforward layers rather than sequential, and untied embedding and unembedding.
@@ G.1 Model Size
G.2 Topic Modeling
The topic model used in Section 5.3 is parameterized as follows: We generate the distribution of the 4 topics in each document using a random Dirichlet distribution with $\alpha=0.3$. The distribution $p(\text{token}\mid\text{topic})$ for each token ${123,132,213,231,312,321}$ is:
@@ G.2 Topic Modeling
Figure 16: Proportion of $S_{3}$ algorithms learned by models first pretrained on the two topic models compared to the randomly initialized baselines. Topic modeling, regardless of the specific distributions, pushs model to learn AA.
G.3 Parity Loss Curriculum
We explored an additional procedure that encourages models to learn parity via an extra loss term. We train an extra linear classifier that takes in the residual activations of an early layer (e.g. layer 3) in which we find parity to be typically computed through probing. We train with this additional loss term to induce the representation to linearly encode parity early on. The classifier is trained to output (1) parity or (2) a (parity, action) tuple, to ensure that the residual also encodes the original action, and not just the parity. After training, we evaluate the model on the original $S_{3}$ task. Both procedures induce the model to learn PAA consistently, though the (parity, action) classifier typically allows the model to generalize better.
G.4 Length Curriculum
We implemented a curriculum training approach where the model was progressively exposed to documents of increasing length. We first trained on only the initial 10 tokens, then expanded to 25 tokens, 50 tokens, and finally the complete 100-token sequences. Each stage of the curriculum was trained for a fixed number of epochs (data). The goal of such a curriculum is to push the model to learn the associative algorithm (AA) as the parity heuristic, we hypothesize, might be less useful for shorter sequences. However, empirically, such a curriculum has no obvious effect on the kind of algorithm the model learns. Of the 5 trials, 2 trials of GPT-2 learn AA, and the other 3 trials learn PAA.
We discover that all five models can perfectly generalize both parity and state when trained on the first 25 tokens. It is only when trained to generalize to sequence length 50 that the distinction between PAA and AA emerges. Models learning the PAA algorithm do not generalize, while models learning AA can generalize from length 25 to 50, as illustrated by Figure 17. This further confirms our finding that the model learns the $S_{3}$ algorithm early on.
Figure 17: Model generalization curves when trained using a length curriculum after training on 10, 25, 50, and 100 tokens respectively. While the length curriculum doesn’t push the model to learn one algorithm or the other, it shows that the model learns these algorithms early on, as indicated by whether the model can generalize well transitioning from training on 25 tokens to 50 tokens.

### v3

@@ Appendix C Full Probing Results
C.1 Probing signatures are relatively consistent across examples
Layer Pythia on S3 (PAA) Pythia on S3 (AA)
State Probe Parity Probe State Probe Parity Probe
0 $4.16\times 10^{-6}$ $3.03\times 10^{-6}$ $6.33\times 10^{-6}$ $2.71\times 10^{-6}$
1 $36\times 10^{-6} 3.3$$31\times 10^{-6}$ $2.62\times 10^{-6}$ $2.82\times 10^{-6}$
2 $4.52\times 10^{-6}$ $5.45\times 10^{-6}$ $1.46\times 10^{-6}$ $8.97\times 10^{-7}$
3 $4.31\times 10^{-5}$ $5.59\times 10^{-4}$ $5.00\times 10^{-6}$ $1.38\times 10^{-6}$
4 $1.58\times 10^{-5}$ $3.72\times 10^{-5}$ $4.16\times 10^{-6}$ $1.11\times 10^{-6}$
5 $1.70\times 10^{-5}$ $5.38\times 10^{-6}$ $97\times 10^{-6} 1.1$$11\times 10^{-6}$
6 $1.87\times 10^{-5}$ $7.02\times 10^{-6}$ $4.73\times 10^{-6}$ $1.59\times 10^{-6}$
7 $1.29\times 10^{-5}$ $1.34\times 10^{-5}$ $7.79\times 10^{-6}$ $3.05\times 10^{-6}$
8 $3.18\times 10^{-5}$ $3.08\times 10^{-5}$ $1.18\times 10^{-5}$ $2.64\times 10^{-6}$
9 $2.96\times 10^{-4}$ $6.85\times 10^{-5}$ $2.36\times 10^{-5}$ $3.53\times 10^{-6}$
10 $3.10\times 10^{-4}$ $7.98\times 10^{-5}$ $2.60\times 10^{-5}$ $2.96\times 10^{-6}$
11 $1.86\times 10^{-4}$ $22\times 10^{-5} 8.8$$83\times 10^{-6}$ $2.77\times 10^{-6}$
12 $7.05\times 10^{-5}$ $4.01\times 10^{-5}$ $1.20\times 10^{-6}$ $2.75\times 10^{-6}$
Table 1: Standard deviations across probe accuracies. We focus our analysis on S3 Pythia models and train 10 probes on different subsets of the S3 dataset. Standard deviations are tiny in all cases, indicating robust signatures.
We investigate the sensitivity of our probing signatures to the data on which the probe was trained. We focus on Pythia models trained on S3. For each model type (PAA vs. AA) and each probe type (parity vs. state probe), we train 10 different probes across 10 different random subsets of the S3 dataset, and report the standard deviations of their accuracies across the 10 runs. Results can be found in Table 1. We find low standard deviations (less than $10^{-3}$) in all four cases, indicating that our probe signatures are robust to different subsets of the data and to randomness in probe training.
C.2 Probe accuracies over sequence lengths
@@
The mean probability the probe put on the correct answer is plotted in Figure 10. We find that, generally speaking, both AA and PAA linearly encode states of exponentially longer sequences as they go down the layers. We find evidence that the PAA models use their intermediate layers to compute parity in parallel: at around the second residual layer, PAA models place $\frac{1}{3}$ probability on the correct answer (there are three actions of each parity in $S_{3}$).
C.3 Examples of Associative Algorithm Representations that Do or Do Not Linearly Encode Parity
@@
Figure 11: Example activations from an AA model that does not linearly encodes parity at the final layer, projected on three principal components with a total explained variance of 41.8%. Blue points have even parity, while orange points have odd parity.
Figure 12: Example activations from an AA model that does linearly encode parity at the final layer. Blue points have even parity, while orange points have odd parity. (Left) State and state parity probe signatures of this model. (Right) projection of hidden representations onto three components with a total explained variance of 49.9%.
@@ D.1 $S_{3}$
We visualize the linear decomposition of the last-layer or penultimate-layer representations across various PAA models. We find the triangular prism shape similar to Figure 4 in all of them, but there was no consistency in which states were paired to form the clusters.
One interpretation is that PAA models may be learning various presentations of $S_{3}$, with each clustering configuration corresponding to a different presentation. Generally speaking, $S_{n}$ can be generated by a 2-cycle and an $n$-cycle: any permutation of $S_{n}$ can be created by composing these two permutations. For example, $S_{3}$ can be generated by the 2-cycle $1\leftrightarrow 2$ and 3-cycle $1\to 2\to 3\to 1$, which corresponds to the clustering $\{(123,213),(312,132),(231,321)\}$: the states within the cluster can be transformed into each other by applying $1\leftrightarrow 2$, while states between clusters are related to each other by $1\to 2\to 3\to 1$. PAA models that cluster according to this pattern may have learned these generators.
@@ D.2 $S_{5}$
What happens in models that learn PAA in $S_{5}$? We visualize the penultimate-layer representation of a Pythia-160M model that learned PAA on $S_{5}$ in 3D space, with parity along one axis and two orthogonal directions along the other two. We find 4 distinct clusters, corresponding to the position of 1 in the state (states having 1 in position 4 and 1 in position 5 are clustered together).
@@
Appendix E Simulating State Tracking in Natural Language
To emulate a more practical scenario, we train pre-trained and from-scratch Pythia models on a version of the S3 permutation composition task expressed in natural language. For example, the permutation “132” would be expressed as “swap positions 2 and 3,” while “312” would be “rotate the last item to the front.” We train LLMs to predict the final state (e.g., 231) from the final period token of the sequence. For example, the following sequence:
Swap positions 2 and 3. Rotate the last item to the front. Swap positions 1 and 2.
would map to permutation sequence “132, 312, 213” and finally correspond to the state “123” after the swaps. We then conduct a similar style of probing and activation patching experiments on models trained on this task.
E.1 Probing experiments
We train probes to map from the activation of each layer at the position of the final “.” token to the final state. Shown in Figure 14, as in our results on synthetic data, probing results are consistent with associative mechanisms – the state probe improves exponentially over layers. In pre-trained Pythia, we see the model represent the state parity much earlier (in terms of layer) than the actual state representation, a signature of the PAA algorithm. For non-pre-trained Pythia, the accuracy of the state and state parity probes increases at a similar rate, indicating that it is more likely learning an associative algorithm; whether it is the strict AA algorithm we identified in the synthetic case is unclear.
E.2 Activation patching experiments
We patch prefixes up to a fixed token position. [^8] Note that token positions may no longer be aligned in natural language: “swap” actions have 5 tokens, and “rotate” actions have 7. Thus, we may not be replacing the activations of the same number of actions between prompts. Nonetheless, we still believe the activation patching results serve as a good proxy for estimating how information gets propagated through the layers. As shown in Figure 14, both models display a distinctly associative signature with exponentially longer prefixes being disregarded for the final state prediction as depth increases. Furthermore, the pre-trained Pythia model possesses a light blue middle section – a sign of the PAA algorithm. Interestingly, the pre-trained Pythia results are significantly more “compressed” over the layers – the LLM computes the state very early on. We suspect this may be due to the pre-trained LLM taking advantage of its innate natural language understanding (and perhaps pre-trained state tracking abilities!) to quickly solve the task in an early layer.
Figure 14: Patching and probing result for Pythia models trained on natural language permutation composition task. We plot the signature of a pre-trained Pythia model (top) and a non-pre-trained Pythia model (bottom). In both cases, the signatures are consistent with the state being learned associatively (both the patching signature and state probe have an exponential curve). The signature of the pre-trained model is consistent with a PAA signature, with the parity probe converging in early layers, and the activation patching signature containing a light-blue middle section.
Appendix F Full Attention Heads Analysis
F.1 Formalizing Parity Heads
@@
Definition F.1.
Let $\alpha_{i,\ell}^{(H)}(x)$ be the attention weight of the $H$th attention head in layer $\ell$ at position $i$ for input $x$ where $x$ is the list of actions $a_{1}\ldots a_{t}]$.
@@
Here, $a_{1}\ldots a_{i}]$ denotes the first $i$ elements of the sequence $x$, and $\text{CI}_{\ell,H,\text{odd}}$ refers to the confidence interval around the average attention weights on odd-parity tokens. We use sequence lengths of up to $L=80$ for $S_{3}$ and $L=50$ for $S_{5}$.
@@
F.2 AA Attention Patterns
Figure 16: Attention patterns in AA models form a tree-like pattern, with tokens in successive layers attending to larger windows of downstream tokens. This is in line with how we expect the associative algorithm to function. We plot only the most salient attention weights by pruning edges for which the attention weight is <0.95<0.95, the target token is not in the top-3 attended-to tokens from the source token, or the source token is not in the top-10 attended-from tokens for the target token. We expect that the attention patterns visualized here do not form a single clean tree, but the superimposition of multiple trees.
@@ F.2 AA Attention Patterns
    Edges between two nodes $(t_{1},l-1)$ and $(t_{2},l)$ if position $t_{2}$ at layer $l$ attends to position $t$ at layer l1l-1. We define “attends to” as follows:
    Let $\alpha_{t_{1}\rightarrow t_{2}}^{(l)}$ denote the maximum attention weight (across all attention heads at layer $l$) from position $t_{1}$ at layer l1l-1 to position $t_{2}$ at layer $l$.
@@
      $\alpha_{t_{1}\rightarrow t_{2}}^{(l)}\in\text{top-10}(\{\alpha_{t_{1}\rightarrow j}^{(l)}:k\in[1,n]\})$: $t_{2}$ is among the top-10 attended-from tokens for $t_{1}$ at layer l1l-1
@@
Note that the tree is not entirely clean: there are redundant edges and edges that cross over each other. We suspect that the subgraph we’ve picked up on is not a single tree, but rather the superimposition of multiple trees, each potentially contributing to not just the prediction for the final token, but also the predictions of the previous tokens.
Appendix G How are these algorithms learned over the course of training?
@@
Appendix H Additional Factors Influencing Learned Algorithm
H.1 Model Size
We have investigated whether model size influences which algorithm the models learn to implement. As shown in Figure 17, model size empirically does not seem to have much effect on the choice of learned algorithm, with model architecture and initialization having a much bigger effect. Some notable differences between the GPT-2 and Pythia architecture are the use of rotary embeddings, parallelized attention, and feedforward layers rather than sequential, and untied embedding and unembedding.
@@ H.1 Model Size
H.2 Topic Modeling
The topic model used in Section 5.3 is parameterized as follows: We generate the distribution of the 4 topics in each document using a random Dirichlet distribution with $\alpha=0.3$. The distribution $p(\text{token}\mid\text{topic})$ for each token 123,132,213,231,312,321{123,132,213,231,312,321} is:
@@ H.2 Topic Modeling
Figure 18: Proportion of $S_{3}$ algorithms learned by models first pre-trained on the two topic models compared to the randomly initialized baselines. Topic modeling, regardless of the specific distributions, pushes the model to learn AA.
H.3 Parity Loss Curriculum
We explored an additional procedure that encourages models to learn parity via an extra loss term. We train an extra linear classifier that takes in the residual activations of an early layer (e.g., layer 3) in which we find parity to be typically computed through probing. We train with this additional loss term to induce the representation to linearly encode parity early on. The classifier is trained to output (1) parity or (2) a (parity, action) tuple, to ensure that the residual also encodes the original action, and not just the parity. After training, we evaluate the model on the original $S_{3}$ task. Both procedures induce the model to learn PAA consistently, though the (parity, action) classifier typically allows the model to generalize better.
H.4 Length Curriculum
We implemented a curriculum training approach where the model was progressively exposed to documents of increasing length. We first trained on only the initial 10 tokens, then expanded to 25 tokens, 50 tokens, and finally the complete 100-token sequences. Each stage of the curriculum was trained for a fixed number of epochs (data). The goal of such a curriculum is to push the model to learn the associative algorithm (AA), as the parity heuristic, we hypothesize, might be less useful for shorter sequences. However, empirically, such a curriculum has no obvious effect on the kind of algorithm the model learns. Of the 5 trials, 2 trials of GPT-2 learn AA, and the other 3 trials learn PAA.
We discovered that all five models can perfectly generalize both parity and state when trained on the first 25 tokens. It is only when trained to generalize to sequence length 50 that the distinction between PAA and AA emerges. Models learning the PAA algorithm do not generalize, while models learning AA can generalize from length 25 to 50, as illustrated by Figure 19. This further confirms our finding that the model learns the $S_{3}$ algorithm early on.
Figure 19: Model generalization curves when trained using a length curriculum after training on 10, 25, 50, and 100 tokens, respectively. While the length curriculum doesn’t push the model to learn one algorithm or the other, it shows that the model learns these algorithms early on, as indicated by whether the model can generalize well from training on 25 tokens to 50 tokens.